### v2

Abstract
For a graph $G$, let $S_{2}(G)$ be the sum of the first two largest signless Laplacian eigenvalues of $G$, and $f(G)=e(G)+3-S_{2}(G)$. Oliveira, Lima, Rama and Carvalho conjectured that $K^{+}_{1,n-1}$ (the star graph with an additional edge) is the unique graph with minimum value of $f(G)$ on $n$ vertices. In this paper, we prove this conjecture, which also confirm a conjecture for the upper bound of $S_{2}(G)$ proposed by Ashraf et al.
Keywords: (Signless) Laplacian matrix; (Signless) Laplacian eigenvalues; Sum of (Signless) Laplacian eigenvalues
1 Introduction
Graphs in this paper are connected and simple. Let $G=(V(G),E(G))$ be a graph, the number of its vertices and edges are denoted by $n(G)$ (or $n$ for short) and $e(G)$, respectively. Let $A(G)$ be the adjacency matrix of $G$ and $D(G)=diag(d(v_{1}),d(v_{2}),\ldots,d(v_{n}))$ be the diagonal matrix of vertex degrees of $G$. The Laplacian matrix and the signless Laplacian matrix of $G$ are defined as $L(G)=D(G)-A(G)$ and $Q(G)=D(G)+A(G)$, respectively. It is well known that both $L(G)$ and $Q(G)$ are positive semidefinite and hence their eigenvalues are nonnegative real numbers. The eigenvalues of $L(G)$ and $Q(G)$ are called Laplacian eigenvalues and signless Laplacian eigenvalues of $G$, respectively, and are denoted by $\mu_{1}(G)\geq\mu_{2}(G)\geq\cdots\geq\mu_{n}(G)$ (or $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{n}$ for short) and $q_{1}(G)\geq q_{2}(G)\geq\cdots\geq q_{n}(G)$ (or $q_{1}\geq q_{2}\geq\cdots\geq q_{n}$ for short), respectively.
In [4], Brouwer proposed the following conjecture for Laplacian eigenvalues.
Conjecture 1.
( [4]) For any graph $G$ with $n$ vertices and for any $k\in\{1,2,\ldots,n\}$,
$\sum_{i=1}^{k}\mu_{i}(G)\leq e(G)+\binom{k+1}{2}.$
It has been proved that Conjecture 1 is true for several classes of graphs such as trees [11], threshold graphs [11], unicyclic graphs [7], [15], bicyclic graphs [7], regular graphs [13] and split graphs [13]. Recently, Cooper [5] showed that Conjecture 1 is true for planar graphs when $k\geq 11$ and for bipartite graphs when $k\geq\sqrt{32}n$. For $k=2$, Haemers et al.[11] showed that Conjecture 1 is true. In [12], Li and Guo proposed the full Brouwer’s Laplacian spectrum conjecture and proved the conjecture holds for $k=2$ with the unique extremal graph.
Let $S_{k}(G)$ be the sum of the first $k$ largest signless Laplacian eigenvalues of $G$. Ashraf et al.[3] proposed the following conjecture.
Conjecture 2.
([3]) For any graph $G$ with $n$ vertices and for any $k\in\{1,2,\ldots,n\}$,
$S_{k}(G)\leq e(G)+\binom{k+1}{2}.$
For bipartite graphs, Conjecture 2 is equivalent to Conjecture 1 since $L(G)$ and $Q(G)$ are similar [6]. Ashraf et al.[3] proved that Conjecture 2 is true for regular graphs, they also proved that Conjecture 2 is true for $k=2$, but the key lemma ([3], Lemma 13) they used is incorrect which has a counterexample in[17]. Zheng[17] proved that Conjecture 2 is true for all connected triangle-free graphs when $k=2$. Therefore, for general graphs Conjecture 2 is still open when $k=2$.
For a graph $G$, define
$f(G)=e(G)+3-S_{2}(G).$
Oliveira et al.[14] conjectured that $K^{+}_{1,n-1}$ (the star graph with an additional edge) minimizes $f(G)$ among all graphs $G$ on $n$ vertices.
Conjecture 3.
([14]) For any graph $G$ on $n\geq 9$ vertices. Then
$f(G)\geq f(K^{+}_{1,n-1}).$
Equality if and only if $G\cong K^{+}_{1,n-1}$.
In this paper, we give a proof of Conjecture 3.
Theorem 1.1.
For any graph $G$ with $n$ vertices, $f(G)\geq f(K^{+}_{1,n-1})$ with equality if and only if $G\cong K^{+}_{1,n-1}$.
By computing the value of $f(K^{+}_{1,n-1})$, one can obtain a tight upper bound of $S_{2}(G)$ which confirm Conjecture 2 in the case $k=2$.
The organization of the remaining of the paper is as follows. In Section 2, we give some lemmas which are useful in the proof of Theorem 1.1. In Section 3, we give the proof of Theorem 1.1.
2 Preliminaries
We begin this section with some definitions and notation. For vertex $v\in V(G)$, the set of its neighbors in $X\subset V(G)$, denoted by $N_{X}(v)$, for convenience, abbreviate $N_{V(G)}(v)$ as $N(v)$. For two graphs $G_{1}$ and $G_{2}$, the union of $G_{1}$ and $G_{2}$, denoted by $G_{1}\cup G_{2}$, is the graph with vertex set $V(G_{1})\cup V(G_{2})$ and edge set $E(G_{1})\cup E(G_{2})$. If $H$ is a subgraph of $G$, let $G-E(H)$ denote the graph obtained from $G$ by removing the edges in $H$.
For any square matrix $M$, we use $P(M,x)$ to denote the characteristic polynomial of $M$. For graph $G$, we write the characteristic polynomial of its signless Laplacian matrix $P(Q(G),x)$ as $P_{Q}(G,x)$.
$H$-join operation of graphs: Let $H$ be a graph with $V(H)=\{1,2,...,k\}$, and $G_{i}$ be disjoint $r_{i}$-regular graphs of order $n_{i}\ (i=1,2,...,k)$. Then the graph $\bigvee_{H}\{G_{1},G_{2},...,G_{k}\}$ is formed by taking the graphs $G_{1},G_{2},...,G_{k}$ and joining every vertex of $G_{i}$ to every vertex of $G_{j}$ whenever $i$ is adjacent to $j$ in $H$. Define $N_{i}=\sum_{j\in N_{H}(i)}n_{j}$ for each $i\in V(H)$ and
$(Q^{\pi})_{ij}=\begin{cases}2r_{i}+N_{i},&i=j,\\ n_{j},&ij\in E(H),\\ 0,&otherwise.\end{cases}$
Lemma 2.1 ([16]).
Let $H$ be a graph with $V(H)=\{1,2,...,k\}$, and $G_{i}$ be a $r_{i}$-regular graph of order $n_{i}\ (i=1,2,...,k)$. If $G=\bigvee_{H}\{G_{1},G_{2},...,G_{k}\}$, then
$P_{Q}(G,x)=P(Q^{\pi},x)\prod_{i=1}^{k}\frac{P_{Q}(G_{i},x-N_{i})}{x-2r_{i}-N_{i}}.$
Lexicographically ordered $k$-sets: Fix a number $k$ between 1 and $n$, inclusive. Consider all possible sets of $k$ distinct numbers between 1 and $n$. Order the elements of each set by using a lexicographic or dictionary order. Denote these sets by $S_{1},S_{2},\ldots,S_{\tbinom{n}{k}}$ so that $S_{1}<S_{2}<\ldots<S_{\tbinom{n}{k}}$.
Compound matrix: Let $A_{i,j}$ be the $k\times k$ matrix formed by the intersection, in $A$, of the rows whose numbers are in the set $S_{i}$ and the columns whose numbers are in the set $S_{j}$. Let $A$ be a complex matrix of order $n$, define $C_{k}(A)$ as its $k$-th compound matrix, which is the matrix of size $\tbinom{n}{k}\times\tbinom{n}{k}$ whose $(i,j)$-th entry is the determinant of the matrix $A_{i,j}$.
Additive compound matrix: The matrix $C_{k}(I+tA)$ is an $\tbinom{n}{k}\times\tbinom{n}{k}$ matrix, each entry of which is a polynomial in $t$ of degree at most $k$. The $k$-th additive compound matrix is defined as
$\Delta_{k}(A)=\frac{d}{dt}C_{k}(I_{n}+tA)\big{|}_{t=0}.$
For additive compound matrix, the following lemma is a well-known result.
Lemma 2.2 ([9]).
If $\lambda_{1},\ \lambda_{2},\ldots,\lambda_{n}$ are all eigenvalues of $A$, then the eigenvalues of $\Delta_{k}(A)$ are the $\tbinom{n}{k}$ distinct sums of the $\lambda_{i}$ taken $k$ at a time.
Corollary 2.1.
If $q_{1},\ q_{2},\ldots,q_{n}$ are signless eigenvalues of $G$, and $P(Q^{\pi},x)$ contains the largest two roots of $P_{Q}(G,x)$, then the largest eigenvalue of $\Delta_{2}(Q^{\pi})$ equals $S_{2}(G)$.
The following result which can be found in matrix theory which plays an important role in our proofs. For a symmetric matrix $A$ of order $n$, let $\lambda_{1}(A)\geq\cdots\geq\lambda_{n}(A)$ be its eigenvalues.
Lemma 2.3 ([8]).
Let $A$ and $B$ be two real symmetric matrices of size $n$. Then for any $1\leq k\leq n$,
$\sum^{k}_{i=1}\lambda_{i}(A+B)\leq\sum^{k}_{i=1}\lambda_{i}(A)+\sum^{k}_{i=1}\lambda_{i}(B).$
Apparently, $\lambda_{1}(\Delta_{k}(A))=\sum^{k}_{i=1}\lambda_{i}(A)$, then the above lemma can be directly obtained from the fact that $\lambda_{1}(\Delta_{k}(A+B))\leq\lambda_{1}(\Delta_{k}(A))+\lambda_{1}(\Delta_{k}(B)).$ An immediate result of Lemma 2.3 is the following corollary which will be used frequently.
Corollary 2.2.
Let $G$ be a graph of order $n$ such that there exist edge disjoint subgraphs of $G$, say $G_{1},\cdots,G_{r}$, with $E(G)=\bigcup^{r}_{i=1}E(G_{i})$. Then $S_{k}(G)\leq\sum^{r}_{i=1}S_{k}(G_{i})$ for any integer $k$ with $1\leq k\leq n$.
The following result is also straightforward.
Lemma 2.4.
For some $k$, if Theorem1.1 holds for $G$ and $H$, then it does for $G\cup H$.
The next lemma is the key to our approach. It gives a sufficient condition for the result of Theorem1.1, that holds for almost all graphs.
Lemma 2.5.
If $G$ is a graph with a nonempty subgraph $H$ for which $S_{k}(H)\leq e(H)$, then $S_{k}(G)<e(G)$.
Proof.
Assume that $G$ is a counterexample with a minimum possible number of edges. By Corollary 2.2, we have $e(G)\leq S_{k}(G)\leq S_{k}(H)+S_{k}(G-H)$. This implies that $S_{k}(G-H)\geq e(G-H)$, since $e(G)\geq e(H)+e(G-H)$ and $S_{k}(H)\leq e(H)$, which contradicts the minimality of $e(G)$. ∎
Corollary 2.3.
If $G$ is a graph with a nonempty subgraph $H$ for which $S_{2}(H)\leq e(H)$, then $S_{2}(G)<e(G)$ and $f(G)>3$.
Lemma 2.6 ([1]).
For any graph $G$ with at most 10 vertices, $f(G)\geq f(K^{+}_{1,n-1})$ with equality if and only if $G\cong K^{+}_{1,n-1}$.
Lemma 2.7 ([6]).
Let $n$ be a positive integer.
  • (1)
    The signless Laplacian eigenvalues of the complete graph $K_{n}$ are $22n-$ 2 and $n-2$ with multiplicities 1 and $n-1$, respectively.
  • (2)
    The signless Laplacian eigenvalues of the star $K_{1,n-1}$ are $n$, 1 and 0 with multiplicities 1, $n-2$, and 1, respectively.
Lemma 2.8 ([10]).
Let $T$ be a tree with $n$ vertices and let $1\leq k\leq n$. Then the sum of the $k$ largest Laplacian eigenvalues of $T$ satisfies $\sum_{i=1}^{k}\mu_{i}(G)\leq(n-1)+2k-1-\frac{2k-2}{n}$. Moreover, equality is achieved only when $k=1$ and $T$ is a star on $n$ vertices.
Using the fact that signless Laplacian matrix and Laplacian matrix are similar for a bipartite graph, from the above lemma, we have the following result.
Corollary 2.4.
Let $G$ be a tree with $n$ vertices, then $f(G)\geq\frac{2}{n}$.
Lemma 2.9 ([6]).
Let $G$ be a graph with $n$ vertices and let $G^{\prime}$ be a graph obtained from $G$ by inserting a new edge into $G$. Then the signless Laplacian eigenvalues of $G$ and $G^{\prime}$ interlace, that is,
$q_{1}(G^{\prime})\geq q_{1}(G)\geq\cdots\geq q_{n}(G^{\prime})\geq q_{n}(G).$
Lemma 2.10.
Let $G$ be a graph with $n$ vertices. Suppose that there exist two non-adjacent vertices $u,v\in V(G)$ such that $q_{k}(G)\geq d(u)+d(v)+2$ for some integer $k$, $1\leq k\leq n$. If $G^{\prime}$ is the graph obtained from $G$ by inserting an edge $e=uv$ into $G$, then $S_{k}(G^{\prime})<S_{k}(G)+1$.
Proof.
For $i=1,\ldots,n$, define $\sigma_{i}=q_{i}(G^{\prime})-q_{i}(G)$. By Lemma 2.9, $\sigma_{i}\geq 0$, for any $i$. Let $d_{1}\geq\cdots\geq d_{n}$ and $d_{1}^{\prime}\geq\cdots\geq d_{n}^{\prime}$ be vertex degrees of $G$ and $G^{\prime}$, respectively. Recall that for any graph $H$, considering the trace of the matrix $Q^{2}(H)$, we have
$\sum^{|V(H)|}_{i=1}q^{2}_{i}(H)=\sum_{v\in V(H)}d^{2}(v)+2e(H).$
Applying this fact, we have
$\sum^{n}_{i=1}q^{2}_{i}(G^{\prime})$ $=\sum^{n}_{i=1}d_{i}^{\prime 2}+2e(G^{\prime})$
$=\sum^{n}_{i=1}d_{i}^{2}+2e(G)+2d(u)+2d(v)+4$
$=\sum^{n}_{i=1}q^{2}_{i}(G)+2(d(u)+d(v)+2).$
This yields that
$2q_{k}(G)\sum^{k}_{i=1}\sigma_{i}$ $\leq\sum^{k}_{i=1}2\sigma_{i}q_{i}(G)$
$<\sum^{n}_{i=1}q^{2}_{i}(G^{\prime})-\sum^{n}_{i=1}q^{2}_{i}(G)$
$=2(d(u)+d(v)+2).$
Since $q_{k}(G)\geq d(u)+d(v)+2$, $S_{k}(G^{\prime})-S_{k}(G)=\sum^{k}_{i=1}\sigma_{i}<1$, and the result follows. ∎
In particular, if $S_{k}(G)<e(G)+\binom{k+1}{2}$, then $S_{k}(G^{\prime})<e(G^{\prime})+\binom{k+1}{2}$.
3 Proof of Theorem1.1
Amaro et al.[1] pointed out that Theorem 1.1 holds for all graphs with at most 10 vertices by computational checking. Assuming the graph we are considering has at least 11 vertices in the next. We firstly present Lemma 3.1 which indicates the graph $K^{+}_{1,n-1}$ is the candidate to be extremal to $f(G)$.
Lemma 3.1.
Let $G$ be isomorphic to $K^{+}_{1,n-1}$. Then $\frac{1.3}{n}<f(G)<\frac{1.5}{n}$.
Proof.
The graph $G$ has an obvious partition $G=\bigvee_{H}\{G_{1},G_{2},G_{3}\}$, and the corresponding quotient matrix is
$Q^{\pi}=\left(\begin{array}[]{ccc}3&1&0\\ 2&n-1&n-3\\ 0&1&1\end{array}\right).$
By Lemma 2.1 and direct calculation, the characteristic polynomial of $Q(G)$ is
$P_{Q}(G,x)=P(Q^{\pi},x)(x-1)^{n-3}.$
The 2-th additive compound matrix of $Q^{\pi}$ is
$\Delta_{2}(Q^{\pi})=\left(\begin{array}[]{ccc}n+2&n-3&0\\ 1&4&1\\ 0&2&n\end{array}\right),$
and its characteristic polynomia is $P(\Delta_{2}(Q^{\pi}),x)=x^{3}-(6+2n)x^{2}+(n^{2}+9n+9)x-3n^{2}-9n+4$. By direct calculation we obtain
$P(\Delta_{2}(Q^{\pi}),n+3-\frac{1.3}{n})$ $>0,$
$P(\Delta_{2}(Q^{\pi}),n+3-\frac{1.5}{n})$ $<0,$
$P(\Delta_{2}(Q^{\pi}),3)$ $>0,$
$P(\Delta_{2}(Q^{\pi}),0)$ $<0.$
Since $q_{2}\geq 2$, by Lemma 2.9, $P(Q^{\pi},x)$ contains the largest two roots of $P_{Q}(G,x)$. From Corollary 2.1 and the fact that $e(G)=n$, we have $\frac{1.3}{n}<f(G)<\frac{1.5}{n}$. ∎
By Lemma 2.4, we may assume that $G$ is a connected graph with at least 11 vertices. Noting that for $H=4K_{2}$ or $H=3K_{1,2}$, one has $S_{2}(H)=e(H)$, using Corollary 2.3, we may assume $G$ contains neither $H=4K_{2}$ nor $H=3K_{1,2}$ as a subgraph. It is sufficient to consider graphs $G$ whose matching number $m(G)$ is at most 3. In the following, we prove Theorem 1.1 for $m(G)=1,2$ and 3, respectively.
Lemma 3.2.
Let $G$ be a graph with $m(G)=1$. Then $f(G)>\frac{1.5}{n}$.
Proof.
Let $n=|V(G)|$. Since $m(G)=1$, it is easy to check that either $G=K_{1,k-1}\cup(n-k)K_{1}$ for some $k$, $1\leq k\leq n$ or $G=K_{3}\cup(n-3)K_{1}$. By Lemma 2.7, the assertion holds. ∎
Let $X,\ Y\subset V(G)$, and $E(X,Y)$ ($e(X,Y)$) be the set of edges (the number of edges) whose one endpoint is in $X$ and one endpoint is in $Y$. A pendant neighbor of $G$ is a vertex adjacent to a pendant vertex. Let $p(G)$ be the number of pendant vertices of $G$ and $q(G)$ be the number of pendant neighbors.
Lemma 3.3.
Let $G$ be a graph with $m(G)=2$. Then $f(G)\geq f(K^{+}_{1,n-1})$ with equality if and only if $G\cong K^{+}_{1,n-1}$.
Proof.
We may assume that $G$ is a connected graph with at least 11 vertices. First suppose that $G$ has no $K_{3}$ as a subgraph, if $G$ is a tree, by Corollary 2.4, we have $f(G)>\frac{1.5}{n}$. Since $m(G)=2$, $G$ is isomorphic to $G_{5}$ in Fig. 1, using the 2-th additive compound matrix approach, we have $f(G)>0.5>\frac{1.5}{n}$, details are deferred in the Appendix.
Next suppose that $G$ has a subgraph $H=K_{3}$ with $V(H)=\{u,v,w\}$.
[…]
If there exists an edge $e=ab$ has no endpoint in $V(H)$, let $M=V(G)-\{a,b,u,v,w\}$, since $m(G)=2$, $M$ is an independent set and $e(V(H),M)=0$, and all vertices in $M$ are adjacent to one of the endpoints of $e$, say $a$. Hence there are no edges between $b$ and $V(H)$, we only need to consider $|N(a)\cap V(H)|=1$ and 3, respectively (isomorphic to $G_{2}$ when $|N(a)\cap V(H)|=2$). Then $G$ is isomorphic to $G_{3}\ (7\leq t=n-4)$ or $G_{4}\ (7\leq t=n-4)$ in Fig. 1.
Figure 1: Possible forms of graphs $G$ with $m(G)=2$
When $G\cong G_{1}$, let $V(\overline{K_{t_{i}}})=\{v_{1},\ldots,v_{t_{i}}\}$ $(i=1,2,3)$, and $G_{t_{1},t_{2},t_{3}}$ be a graph of order $n$ $(n=t_{1}+t_{2}+t_{3})$, which consists of three independent sets $\overline{K_{t_{i}}}$ $(i=1,2,3)$, where $t_{1}$ triangles share one common edge and $t_{2},t_{3}$ pendant vertices at each end of the common edge. The graph $G_{t_{1},t_{2},t_{3}}$ has an obvious partition $G=\bigvee_{H}\{G_{1},G_{2},...,G_{k}\}(k=5)$ where the corresponding quotient matrix is
$Q^{\pi}=\left(\begin{array}[]{ccccc}2&0&0&1&1\\ 0&1&0&1&0\\ 0&0&1&0&1\\ t_{1}&t_{2}&0&t_{1}+t_{2}+1&1\\ t_{1}&0&t_{3}&1&t_{1}+t_{3}+1\end{array}\right).$
By Lemma 2.1 and direct calculation, the characteristic polynomial of $Q(G_{t_{1},t_{2},t_{3}})$ is
$P_{Q}(G_{t_{1},t_{2},t_{3}},x)=P(Q^{\pi}(G_{t_{1},t_{2},t_{3}}),x)(x-2)^{t_{1}-1}(x-1)^{t_{2}+t_{3}-2}.$
Without loss of generality, we may assume $t_{2}\geq t_{3}$. Let
$P_{1}$ $=P(\Delta_{2}(Q^{\pi}(G_{t_{1},t_{2},t_{3}})),x+e(G_{t_{1},t_{2},t_{3}})+3),$
$P_{2}$ $=P(\Delta_{2}(Q^{\pi}(G_{t_{1},t_{2}+1,t_{3}-1})),x+e(G_{t_{1},t_{2}+1,t_{3}-1})+3).$
By direct computation, $P_{1}-P_{2}$ is always positive (details are deferred in the Appendix), which indicates the largest root of $P_{1}$ less than the largest root of $P_{2}$.
Since $q_{2}>2$, by Lemma 2.9 (if $t_{1}=1$ and $t_{3}=0$, the previous results (see Theorem 2.4 in [2]) verify that $q_{2}>3-\frac{2.5}{n}>2$), $P(Q^{\pi}(G_{t_{1},t_{2},t_{3}},x))$ and $P(Q^{\pi}(G_{t_{1},t_{2}+1,t_{3}-1},x))$ contain the largest two roots of $P_{Q}(G_{t_{1},t_{2},t_{3}},x)$ and $P_{Q}(G_{t_{1},t_{2}+1,t_{3}-1},x)$. By Corollary 2.1, we have $f(G_{t_{1},t_{2},t_{3}})>f(G_{t_{1},t_{2}+1,t_{3}-1})$. By recursive relationships, we have $f(G_{1})>f(G_{1}^{1})$.
Let $G_{t_{1},t_{2}}$ be a graph of order $n$ $(n=t_{1}+t_{2})$, which consists of two independent sets $\overline{K_{t_{i}}}$ $(i=1,2)$, where $t_{1}$ triangles share one common edge and $t_{2}$ pendant vertices at the end of the common edge. Let
$P_{3}$ $=P(\Delta_{2}(Q^{\pi}(G_{t_{1},t_{2}})),x+e(G_{t_{1},t_{2}})+3),$
$P_{4}$ $=P(\Delta_{2}(Q^{\pi}(G_{t_{1}-1,t_{2}+1})),x+e(G_{t_{1}-1,t_{2}+1})+3).$
By using computer computations, $P_{3}-P_{4}$ is always positive (details are deferred in the Appendix), which indicates the largest root of $P_{3}$ less than the largest root of $P_{4}$.
Since $q_{2}>2$, by Lemma 2.9 (if $t_{1}=1$, the previous results (see Theorem 2.4 in [2]) verify that $q_{2}>3-\frac{2.5}{n}>2$), $P(Q^{\pi}(G_{t_{1},t_{2}},x))$ and $P(Q^{\pi}(G_{t_{1}-1,t_{2}+1},x))$ contains the largest two roots of $P_{Q}(G_{t_{1},t_{2}},x)$ and $P_{Q}(G_{t_{1}-1,t_{2}+1},x)$. By Corollary 2.1, we have $f(G_{t_{1},t_{2}})>f(G_{t_{1}-1,t_{2}+1})$. By recursive relationships, we have $f(G_{1}^{1})>f(G_{1}^{2})$. Then $f(G_{1})>f(G_{1}^{1})>f(G_{1}^{2})=f(K^{+}_{1,n-1})$.
When $G\cong G_{2}$, the graph $G$ has an obvious partition $G=\bigvee_{H}\{G_{1},G_{2},...,G_{k}\}(k=4)$ where the corresponding quotient matrix is
$Q^{\pi}=\left(\begin{array}[]{cccc}1&1&0&0\\ t&t+2&2&0\\ 0&1&4&1\\ 0&0&2&2\end{array}\right).$
By Lemma 2.1 and direct calculation, the characteristic polynomial of $Q(G)$ is
$P_{Q}(G,x)=P(Q^{\pi},x)(x-2)(x-1)^{t-1}.$
By simple calculations we obtain
$P(Q^{\pi},t+3.25-\frac{1.5}{t+4})$ $=\frac{0.25}{(4+t)^{4}}g(t)>0,$
$P(Q^{\pi},4.75)$ $=-0.3t-19.98<0,$
$P(Q^{\pi},2)$ $=4t>0,$
$P(Q^{\pi},1)$ $=-t<0,$
$P(Q^{\pi},0)$ $=8>0,$
with $g(t)=(t^{2}+5.3t+8.9)(t+5.2)(t+4.6)(t+4.3)(t+1.3)(t-6.9)$, which indicates $q_{1}(G)<t+3.25-\frac{1.5}{t+4}$ and $q_{2}(G)<4.75$, then $S_{2}(G)<t+8-\frac{1.5}{t+4}=e(G)+3-\frac{1.5}{n}$, we have $f(G)>\frac{1.5}{n}$.
When $G\cong G_{3}$, we use the quotient matrix approach, when $G\cong G_{4}$, we use the 2-th additive compound matrix approach, details are deferred in the Appendix. This completes the proof.
For nonempty set $X\subset V(G)$, let $N(X)=\{v\mid v\not\in X\ \text{and}\ N(v)\cap X\neq\emptyset\}$. Suppose that $V^{\prime}$, $E^{\prime}$ are nonempty subsets of $V(G)$ and $E(G)$, the subgraphs induced by $V^{\prime}$ and $E^{\prime}$ are denoted by $G[V^{\prime}]$ and $G[E^{\prime}]$, respectively.
Lemma 3.4.
Let $G$ be a graph with $m(G)=3$. Then $f(G)>\frac{1.5}{n}$.
Proof.
Without loss of generality, we may suppose that $e_{1}=a_{1}b_{1}$, $e_{2}=a_{2}b_{2}$ and $e_{3}=a_{3}b_{3}$ are three independent edges in $G$.
We first suppose that $G$ contains $K_{3}\cup 2K_{2}$ as a subgraph and $V(K_{3})=\{a_{1},b_{1},u\}$. Let $V_{1}=V(K_{3}\cup 2K_{2})$, $M^{\prime}=V(G)-V_{1}$. Then $|M^{\prime}|\geq 4$. Since $m(G)=3$, $M^{\prime}$ is an independent set and $e(V(K_{3}),M^{\prime})=0$. Since $G$ has neither $3K_{1,2}$ nor $4K_{2}$ as a subgraph and $|M^{\prime}|\geq 4$, there exists exactly one vertex $a_{3}$ (w.l.o.g.) which is adjacent to vertices of $M^{\prime}$ and $d(b_{3})=1$.
Let $V_{2}=\{a_{1},b_{1},u,a_{2},b_{2}\}$, $E_{1}=\{a_{3}w|w\in V_{2}\}$. Then $G[E_{1}]$ is a star with center $a_{3}$, and $G-E_{1}$ is a disjoint union of a star $K_{1,|M^{\prime}|+1}$ with center $a_{3}$ and a graph $G[V_{2}]$ containing $K_{3}+K_{2}$.
If $G[V_{2}]$ is connected, then $G[V_{2}]$ is spanning subgraph of $K_{5}$ which contains $K_{3}\cup K_{2}$ as a subgraph, by direct calculation, we have $q_{1}(G[V_{2}])<e(G[V_{2}])-0.15<e(G[V_{2}])-\frac{1.5}{n}\ (n\geq 11)$. Since $n(K_{1,|M^{\prime}|+1})\geq 6$, so $q_{2}(G[V_{2}])\leq q_{2}(K_{5})<q_{1}(K_{1,|M^{\prime}|+1})$. Hence $S_{2}(G-E_{1})=q_{1}(G[V_{2}])+q_{1}(K_{1,|M^{\prime}|+1})<e(G[V_{2}])-\frac{1.5}{n}+e(K_{1,|M^{\prime}|+1})+1=e(G-E_{1})+1-\frac{1.5}{n}$. Thus by Corollary 2.2, $S_{2}(G)\leq S_{2}(G[E_{1}])+S_{2}(G-E_{1})<e(G[E_{1}])+2+e(G-E_{1})+1-\frac{1.5}{n}=e(G)+3-\frac{1.5}{n}$, we have $f(G)>\frac{1.5}{n}$.
If $G[V_{2}]$ is disconnected, then $G[V_{2}]=K_{3}\cup K_{2}$. Let $t=d(a_{2})+d(b_{2})$. We have $t=3,4$. It is not hard to see that $m(G-e_{2})=2$ and $G-e_{2}$ contains disjoint union of a star $K_{1,|M^{\prime}|+1}$ with center $a_{3}$ and $K_{3}$. Therefore, by Lemma 2.9, $q_{2}(G-e_{2})\geq t$. Using Lemmas 3.3 and 2.10, we find that $S_{2}(G)<S_{2}(G-e_{2})+1<(e(G-e_{2})+3-\frac{1.5}{n})+1=e(G)+3-\frac{1.5}{n}$, we have $f(G)>\frac{1.5}{n}$.
Next suppose that $G$ has no subgraph $K_{3}\cup 2K_{2}$. Since $m(G)=3$, $M=V(G)-V(\{e_{1},e_{2},e_{3}\})$ is an independent set and $|M|\geq 5$. Since $G$ has no $4K_{2}$ and $K_{3}+2K_{2}$ as subgraph, either $N_{M}(a_{i})=\emptyset$ or $N_{M}(b_{i})=\emptyset$ for $i=1,2,3$. Without loss of generality, we may assume that $N(M)\subseteq\{a_{1},a_{2},a_{3}\}$. We consider the following three cases.
Case 1: $|N(M)|=3$.
We have $N(M)=\{a_{1},a_{2},a_{3}\}$. The bipartite subgraph $G[E(\{a_{1},a_{2},a_{3}\},M)]$ has no perfect matching since $G$ has no $3K_{1,2}$. By Hall’s Theorem, there exists a subset of $\{a_{1},a_{2},a_{3}\}$ with 2 elements, say $\{a_{2},a_{3}\}$, such that $|N(\{a_{2},a_{3}\})\cap M|=1$ (we have $|N(\{a_{2},a_{3}\})\cap M|\neq 0$ since $|N(M)|=3$). This means that there exists exactly one vertex $y\in M$ which is adjacent to both $a_{2}$ and $a_{3}$. If $d(b_{1})\geq 2$, then we clearly find a subgraph isomorphic to $3K_{1,2}$ in $G$, a contradiction. Therefore, $d(b_{1})=1$. Let $V_{1}=\{a_{2},b_{2},a_{3},b_{3},y\}$, $E_{1}=\{a_{1}w|w\in V_{1}\}$. Then $G[E_{1}]$ is a star with center $a_{1}$, and $G-E_{1}$ is a disjoint union of a star $K_{1,|M|}$ with center $a_{1}$ and a graph $G[V_{1}]$ containing $P_{5}$. Since G has no $K_{3}\cup 2K_{2}$, two edges $a_{2}a_{3}$ and $b_{2}b_{3}$ can not exist at the same time, therefore $G[V_{1}]$ is isomorphic to graphs in Fig. 2.
Figure 2: Possible forms of graphs containing a $P_{5}$
By direct calculation, we have $q_{1}(G[V_{1}])<e(G[V_{1}])-\frac{1.5}{n}\ (n\geq 11)$. Since $n(K_{1,|M|})\geq 6$, so $q_{2}(G[V_{1}])\leq q_{1}(K_{1,|M|})$. Hence $S_{2}(G-E_{1})=q_{1}(G[V_{1}])+q_{1}(K_{1,|M|})<e(G[V_{1}])-\frac{1.5}{n}+e(K_{1,|M|})+1=e(G-E_{1})+1-\frac{1.5}{n}$. Thus by Corollary 2.2, $S_{2}(G)\leq S_{2}(G[E_{1}])+S_{2}(G-E_{1})<e(G[E_{1}])+2+e(G-E_{1})+1-\frac{1.5}{n}=e(G)+3-\frac{1.5}{n}$, we have $f(G)>\frac{1.5}{n}$.
Case 2: $|N(M)|=2$.
Without loss of generality, suppose that $N(M)=\{a_{1},a_{2}\}$. Since $m(G)=3$, $b_{1}$ is not adjacent to $b_{2}$. If $b_{1}$ is adjacent to $a_{3}$ or $b_{3}$, then changing the role of $e_{1}$, $e_{2}$, $e_{3}$ by three independent edges $\{a_{1},z\}$, $e_{2}$, $e_{3}$ for some vertex $z\in M\cap N(a_{1})$, we have Case 1. Therefore, we may assume that $b_{1}$, and similarly $b_{2}$, is adjacent to none of the vertices $a_{3}$ and $b_{3}$. Hence we only need to consider $|N(\{a_{3},b_{3}\})|=1$ or 2, respectively. Consider the existence of the edge $a_{1}a_{2}$, possible forms of graph $G$ are in Fig. 3.
Figure 3: Possible forms of graphs $G$ with $m(G)=3$ and $|N(M)|=2$
When $G\cong G_{i}$, $i=6,7,8,9,10,12,13$, let $t=d(a_{3})+d(b_{3})$, we have $t\in\{3,4,5\}$.
We first present several special cases: $t_{1}=t_{2}=1$ when $G\cong G_{6}$; $t_{1}=0$ and $t_{2}=2$; $t_{1}=t_{2}=1$; $t_{1}=2$ and $t_{2}=0$ when $G\cong G_{8},G_{10},G_{13}$ (were checked by computer computations).
Except these special cases, it is not hard to see that $m(G-e_{3})=2$ and $G-e_{3}$ contains two disjoint stars $K_{1,t-1}$ with centers $a_{1}$ and $a_{2}$ Therefore, by Lemma 2.9, $q_{2}(G-e_{3})\geq t$. Using Lemmas 3.3 and 2.10, we find that $S_{2}(G)<S_{2}(G-e_{3})+1<(e(G-e_{3})+3-\frac{1.5}{n})+1=e(G)+3-\frac{1.5}{n}$, we have $f(G)>\frac{1.5}{n}$.
When $G\cong G_{11}$, we use the 2-th additive compound matrix approach, by using computer computations and recursive relationships, we have $f(G_{11})>f(G_{11}^{1})$ and $f(G_{11}^{1})>f(G_{11}^{2})$. Then $f(G_{11})>f(G_{11}^{1})>f(G_{11}^{2})=f(G_{2})$, we have $f(G)>\frac{1.5}{n}$.
When $G\cong G_{14}$, we use the 2-th additive compound matrix approach, by using computer computations and recursive relationships, we have $f(G_{14})>f(G_{14}^{1})$ and $f(G_{14}^{1})>f(G_{14}^{2})$. Then $f(G_{14})>f(G_{14}^{1})>f(G_{14}^{2})=f(G_{4})$, we have $f(G)>\frac{1.5}{n}$.
Case 3: $|N(M)|=1$.
Without loss of generality, suppose that $N(M)=\{a_{1}\}$. If $d(b_{1})\geq 2$, then we clearly find three independent edges $e^{\prime}_{1}$, $e^{\prime}_{2}$, $e^{\prime}_{3}$ in $G$ such that the set $M^{\prime}=V(G)-V(\{e^{\prime}_{1},e^{\prime}_{2},e^{\prime}_{3}\}$ is an independent set and $|N(M^{\prime})|\geq 2$ which is dealt with as the previous cases. Therefore, $d(b_{1})=1$. Let $V_{1}=\{a_{2},b_{2},a_{3},b_{3}\}$, $E_{1}=\{a_{1}w|w\in V_{1}\}$. Then $G[E_{1}]$ is a star with center $a_{1}$, and $G-E_{1}$ is a disjoint union of a star $K_{1,|M|+1}$ with center $a_{1}$ and a graph $G[V_{1}]$ containing $2K_{2}$. Therefore $G[V_{1}]$ is isomorphic to graphs in Fig. 4.
Figure 4: Possible forms of graphs containing $2K_{2}$
By direct calculation, we have $q_{1}(G[V_{1}])<e(G[V_{1}])+1-\frac{1.5}{n}\ (n\geq 11)$, hence $S_{2}(G-E_{1})=q_{1}(G[V_{1}])+q_{1}(K_{1,|M|+1})<e(G[V_{1}])+1-\frac{1.5}{n}+e(K_{1,|M|})+1=e(G-E_{1})+2-\frac{1.5}{n}$. If $|N(a_{1})\cap V_{1}|=1$, which means $|E_{1}|=1$, by Corollary 2.2, $S_{2}(G)\leq S_{2}(G[E_{1}])+S_{2}(G-E_{1})<e(G[E_{1}])+1+e(G-E_{1})+2-\frac{1.5}{n}=e(G)+3-\frac{1.5}{n}$, we have $f(G)>\frac{1.5}{n}$. Hence we only need to consider $|N(a_{1})|\cap V_{1}|=2$, 3 and 4, respectively. Possible forms of graph $G$ are in Fig. 5 (If $G$ is a tree graph, the assertion holds by Corollary 2.4).
Figure 5: Possible forms of graphs $G$ with $m(G)=3$ and $|N(M)|=1$
When $G\cong G_{i}$, $i=17,24,25$, let $t=d(a_{3})+d(b_{3})$ (pick $a_{2}$ and $b_{2}$ when $G\cong G_{25}$), we have $t\in\{3,4\}$. It is not hard to see that $m(G-e_{3})=2$ and $G-e_{3}$ contains two disjoint stars $K_{1,t-1}$ with centers $a_{1}$ and $b_{2}$ (contains disjoint union of a star $K_{1,t-1}$ with center $a_{1}$ and a $K_{3}$ when $G\cong G_{25}$). Therefore, by Lemma 2.9, $q_{2}(G-e_{3})\geq t$. Using Lemmas 3.3 and 2.10, we find that $S_{2}(G)<S_{2}(G-e_{3})+1<(e(G-e_{3})+3-\frac{1.5}{n})+1=e(G)+3-\frac{1.5}{n}$, we have $f(G)>\frac{1.5}{n}$.
When $G$ is isomorphic to other graphs in Fig. 5, we use the quotient matrix approach, details are deferred in the Appendix. This completes the proof. ∎
Proof of Theorem 1.1: The result follows immediately from Corollary 2.3, Lemmas 3.1, 3.2, 3.3 and 3.4.
4 Declaration of competing interest
There is no competing interest.
Appendix
The details of polynomial $P_{1}-P_{2}$ and $P_{3}-P_{4}$ can be obtained from:
https://github.com/ZZM0329/Polynomial/tree/main/G1
Table 1: 2-th additive compound matrix approach
No. Corresponding quotient matrix $Q^{\pi}$ Characteristic polynomial $P_{Q}(G,x)$
$G_{4}$ $\left(\begin{array}[]{ccc}1&1&0\\ t&t+3&3\\ 0&1&5\end{array}\right)$ $P(Q^{\pi},x)(x-2)^{2}(x-1)^{t-1}$
$G_{5}$ $\left(\begin{array}[]{ccccc}2&0&0&1&1\\ 0&1&0&1&0\\ 0&0&1&0&1\\ t_{1}&t_{2}&0&t_{1}+t_{2}&0\\ t_{1}&0&t_{3}&0&t_{1}+t_{3}\end{array}\right)$ $P(Q^{\pi},x)(x-2)^{t_{1}-1}(x-1)^{t_{2}+t_{3}-2}$
$G_{11}$ $\left(\begin{array}[]{cccccc}2&0&0&1&1&0\\ 0&1&0&1&0&0\\ 0&0&1&0&1&0\\ t_{1}&t_{2}&0&t_{1}+t_{2}+2&0&2\\ t_{1}&0&t_{3}&0&t_{1}+t_{3}+2&2\\ 0&0&0&1&1&4\end{array}\right)$ $P(Q^{\pi},x)(x-2)^{t_{1}}(x-1)^{t_{2}+t_{3}-2}$
$G_{14}$ $\left(\begin{array}[]{cccccc}2&0&0&1&1&0\\ 0&1&0&1&0&0\\ 0&0&1&0&1&0\\ t_{1}&t_{2}&0&t_{1}+t_{2}+3&1&2\\ t_{1}&0&t_{3}&1&t_{1}+t_{3}+3&2\\ 0&0&0&1&1&4\end{array}\right)$ $P(Q^{\pi},x)(x-2)^{t_{1}}(x-1)^{t_{2}+t_{3}-2}$
Table 2: Quotient matrix approach
No. Corresponding quotient matrix $Q^{\pi}$
Characteristic polynomial $P_{Q}(G,x)$
$G_{3}$ $\left(\begin{array}[]{cccc}1&1&0&0\\ t&t+1&1&0\\ 0&1&3&2\\ 0&0&1&3\end{array}\right)$
$[x^{4}-(t+8)x^{3}+(6t+19)x^{2}-(7t+16)x+4](x-1)^{t}$
$G_{15}$ $\left(\begin{array}[]{ccccc}1&1&0&0&0\\ t&t+3&2&0&1\\ 0&1&3&0&0\\ 0&0&0&1&1\\ 0&1&0&1&2\end{array}\right)$
$[x^{5}-(t+10)+x^{4}+(6t+34)x^{3}-(10t+48)x^{2}+(3t+27)x-4](x-1)^{t}$
$G_{16}$ $\left(\begin{array}[]{cccc}1&1&0&0\\ t&t+4&2&2\\ 0&1&3&0\\ 0&1&0&3\end{array}\right)$
$[x^{3}-(t+8)x^{2}+(3t+15)x-8](x-3)(x-1)^{t+1}$
$G_{18}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+2&1&0&1&0\\ 0&1&2&1&0&0\\ 0&0&1&2&1&0\\ 0&1&0&1&3&1\\ 0&0&0&0&1&1\end{array}\right)$
$[x^{4}-(t+10)x^{3}+(7t+34)x^{2}-(13t+46)x+4t+20](x-1)^{t}x$
$G_{19}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+2&1&0&0&1\\ 0&1&2&1&0&0\\ 0&0&1&2&1&0\\ 0&0&0&1&2&1\\ 0&1&0&0&1&2\end{array}\right)$
$[x^{4}-(t+8)x^{3}+(5t+20)x^{2}-(5t+17)x+4](x^{2}-3x+1)(x-1)^{t-1}$
$G_{20}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+2&0&1&1&0\\ 0&0&1&1&0&0\\ 0&1&1&3&1&0\\ 0&1&0&1&3&1\\ 0&0&0&0&1&1\end{array}\right)$
$[x^{4}-(t+8)x^{3}+(5t+18)x^{2}-(3t+15)x+4](x^{2}-3x+1)(x-1)^{t-1}$
$G_{21}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+3&1&1&1&0\\ 0&1&2&1&0&0\\ 0&1&1&3&1&0\\ 0&1&0&1&3&1\\ 0&0&0&0&1&1\end{array}\right)$
$[x^{6}-(t+13)x^{5}+(9t+62)x^{4}-(26t+140)x^{3}+(27t+158)x^{2}-(8t+84)x+16](x-1)^{t-1}$
$G_{22}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+3&1&1&0&1\\ 0&1&2&1&0&0\\ 0&1&1&3&1&0\\ 0&0&0&1&2&1\\ 0&1&0&0&1&2\end{array}\right)$
$[x^{6}-(t+13)x^{5}+(9t+63)x^{4}-(27t+145)x^{3}+(31t+165)x^{2}-(11t+87)x+16](x-1)^{t-1}$
$G_{23}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+4&1&1&1&1\\ 0&1&2&1&0&0\\ 0&1&1&3&1&0\\ 0&1&0&1&3&1\\ 0&1&0&0&1&2\end{array}\right)$
$[x^{4}-(t+11)x^{3}+(6t+37)x^{2}-(7t+47)x+20](x-3)(x-1)^{t}$
$G_{26}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+2&1&0&0&1\\ 0&1&2&1&0&0\\ 0&0&1&3&1&1\\ 0&0&0&1&2&1\\ 0&1&0&1&1&3\end{array}\right)$
$[x^{5}-(t+12)x^{4}+(9t+51)x^{3}-(24t+94)x^{2}+(19t+71)x-16](x-1)^{t}$
$G_{27}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+2&0&1&0&1\\ 0&0&1&1&0&0\\ 0&1&1&4&1&1\\ 0&0&0&1&2&1\\ 0&1&0&1&1&3\end{array}\right)$
$[x^{6}-(t+13)x^{5}+(10t+61)x^{4}-(31t+135)x^{3}+(35t+150)x^{2}-(12t+80)x+16](x-1)^{t-1}$
$G_{28}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+3&1&1&0&1\\ 0&1&2&1&0&0\\ 0&1&1&4&1&1\\ 0&0&0&1&2&1\\ 0&1&0&1&1&3\end{array}\right)$
$[x^{6}-(t+15)x^{5}+(11t+84)x^{4}-(40t+228)x^{3}+(58t+319)x^{2}-(29t+221)x+60](x-1)^{t-1}$
$G_{29}$ $\left(\begin{array}[]{ccccc}1&1&0&0&0\\ t&t+3&1&0&2\\ 0&1&2&1&0\\ 0&0&1&3&2\\ 0&1&0&1&4\end{array}\right)$
$[x^{5}-(t+13)x^{4}+(9t+59)x^{3}-(23t+115)x^{2}+(16t+92)x-24](x-2)(x-1)^{t-1}$
$G_{30}$ $\left(\begin{array}[]{ccccc}1&1&0&0&0\\ t&t+3&0&1&2\\ 0&0&1&1&0\\ 0&1&1&4&2\\ 0&1&0&1&4\end{array}\right)$
$[x^{5}-(t+13)x^{4}+(9t+57)x^{3}-(21t+107)x^{2}+(10t+86)x-24](x-2)(x-1)^{t-1}$
$G_{31}$ $\left(\begin{array}[]{ccccc}1&1&0&0&0\\ t&t+4&1&1&2\\ 0&1&2&1&0\\ 0&1&1&4&2\\ 0&1&0&1&4\end{array}\right)$
$[x^{4}-(t+12)x^{3}+(7t+43)x^{2}-(8t+56)x+24](x-3)(x-2)(x-1)^{t-1}$
$G_{32}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+2&1&1&0&0\\ 0&1&3&1&0&1\\ 0&1&1&3&1&0\\ 0&0&0&1&2&1\\ 0&0&1&0&1&2\end{array}\right)$
$[x^{4}-(t+10)x^{3}+(7t+32)x^{2}-(11t+39)x+16](x^{2}-3x+1)(x-1)^{t-1}$
$G_{33}$ $\left(\begin{array}[]{cccc}1&1&0&0\\ t&t+2&0&2\\ 0&0&2&2\\ 0&1&2&3\end{array}\right)$
$[x^{3}-(t+8)x^{2}+(5t+17)x-2t-10](x-3)(x-2)(x-1)^{t-1}x$
$G_{34}$ $\left(\begin{array}[]{ccccc}1&1&0&0&0\\ t&t+3&1&2&0\\ 0&1&3&2&0\\ 0&1&1&3&1\\ 0&0&0&2&2\end{array}\right)$
$[x^{5}-(t+12)x^{4}+(8t+49)x^{3}-(17t+86)x^{2}+(8t+64)x-16](x-3)(x-1)^{t-1}$
$G_{35}$ $\left(\begin{array}[]{ccc}1&1&0\\ t&t+4&4\\ 0&1&5\end{array}\right)$
$[x^{3}-(t+10)x^{2}+(5t+25)x-16](x-3)^{2}(x-1)^{t}$
$G_{36}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+2&1&1&0&0\\ 0&1&4&1&1&1\\ 0&1&1&3&1&0\\ 0&0&1&1&3&1\\ 0&0&1&0&1&2\end{array}\right)$
$[x^{6}-(t+15)x^{5}+(12t+84)x^{4}-(48t+228)x^{3}+(77t+319)x^{2}-(41t+221)x+60](x-1)^{t-1}$
$G_{37}$ $\left(\begin{array}[]{cccc}1&1&0&0\\ t&t+2&0&2\\ 0&0&4&2\\ 0&1&2&3\end{array}\right)$
$[x^{4}-(t+10)x^{3}+(7t+29)x^{2}-(8t+28)x+8](x-3)(x-2)(x-1)^{t-1}$
$G_{38}$ $\left(\begin{array}[]{cccc}1&1&0&0\\ t&t+2&2&0\\ 0&1&5&2\\ 0&0&2&2\end{array}\right)$
$[x^{3}-(t+9)x^{2}+(6t+18)x-8](x-3)(x-2)(x-1)^{t}$
$G_{39}$ $\left(\begin{array}[]{ccccc}1&1&0&0&0\\ t&t+3&1&2&0\\ 0&1&4&2&1\\ 0&1&1&3&1\\ 0&0&1&2&3\end{array}\right)$
$[x^{5}-(t+14)x^{4}+(10t+68)x^{3}-(28t+146)x^{2}+(23t+139)x-48](x-3)(x-1)^{t-1}$
$G_{40}$ $\left(\begin{array}[]{ccccc}1&1&0&0&0\\ t&t+3&2&1&0\\ 0&1&5&1&1\\ 0&1&2&3&0\\ 0&0&2&0&2\end{array}\right)$
$[x^{5}-(t+14)x^{4}+(10t+67)x^{3}-(27t+142)x^{2}+(20t+136)x-48](x-3)(x-1)^{t-1}$
$G_{41}$ $\left(\begin{array}[]{cccc}1&1&0&0\\ t&t+4&2&2\\ 0&1&5&2\\ 0&1&2&3\end{array}\right)$
$[x^{4}-(t+13)x^{3}+(8t+51)x^{2}-(11t+75)x+36](x-3)^{2}(x-1)^{t-1}$
$G_{42}$ $\left(\begin{array}[]{cccc}1&1&0&0\\ t&t+2&2&0\\ 0&1&5&2\\ 0&0&2&4\end{array}\right)$
$[x^{4}-(t+12)x^{3}+(9t+43)x^{2}-(16t+56)x+24](x-3)(x-2)(x-1)^{t-1}$
$G_{43}$ $\left(\begin{array}[]{ccccc}1&1&0&0&0\\ t&t+3&2&1&0\\ 0&1&5&1&1\\ 0&1&2&4&1\\ 0&0&2&1&3\end{array}\right)$
$[x^{4}-(t+13)x^{3}+(9t+51)x^{2}-(15t+75)x+36](x-3)^{2}(x-1)^{t-1}$
$G_{44}$ $\left(\begin{array}[]{ccc}1&1&0\\ t&t+4&4\\ 0&1&7\end{array}\right)$
$[x^{3}-(t+12)x^{2}+(7t+35)x-24](x-3)^{3}(x-1)^{t-1}$